OpenStudy (anonymous):

I need help . 12. A certain object has an initial height of 40 feet and a maximum height of 500 feet. Find the height equation, s(t) = -16t2 + v0t + s0, for the object. (Hint: With a vertex of (h,k), the height equation can be written as s(t) = -16(t -h)² + k.

OpenStudy (anonymous):

 y-coord. = 500 s(t) = -16(t -h)² + 500 initial height is 40 therefore, s(0)=40 (intial i.e. t=0) -16(0-h)² + 500= 40 -16h^2 = -460 h^2 = 115/4 h=5.36 therefore, the equation is s(t) = -16(t -(sqrt(115/4)))² + 500 then expand it.